OpenStudy (anonymous):

find the measure of an angle if its measure is 5 degrees less than four times the measure of its complement

OpenStudy (anonymous):

angle + its complement = 90 degree Let the complement angle = x degrees then the angle = (4x - 5) degrees Since their sum has to be 90 degree x + (4x - 5) = 90 Now solve to get value of x...

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

x + 4x - 5 = 90 5x - 5 = 90 5x = 90 - 5 x = 85/5 x = 17

OpenStudy (anonymous):

The answer choices are 19 degrees, 37, 143, and 71 though.

OpenStudy (anonymous):

I made a slight mistake, here is the correct solution x + 4x - 5 = 90 5x - 5 = 90 5x = 90 + 5 x = 95/5 x = 19 The required angle is 4x - 5 = 4(19) - 5 = 76 - 5 = 71

OpenStudy (anonymous):

Thank you!